OpenStudy (anonymous):

find the tangents of the curve 2x^2 + y^2 = 54 from (10, 1) please help. the answers are x-y -9 = 0 and 53x + 73y - 603= 0

OpenStudy (anonymous):

its just getting the dy/dx

OpenStudy (wolfe8):

Getting the dy/dx will give you the gradient of the tangent to the line at a given point. What else would you need for a tangent(straight line) equation? Also, why are there 2 answers?

OpenStudy (anonymous):

2 slopes

OpenStudy (wolfe8):

Ah I get it. So what is the dy/dx you get? And what does it represent?

OpenStudy (anonymous):

the slope..

OpenStudy (anonymous):

i got -4x/2y

OpenStudy (anonymous):

@phi

OpenStudy (wolfe8):

First you rearrange the equation to solve for y: \[y=\pm \sqrt{54-2x ^{2}}\]I would assume you will have 2 answers because you will get 2 slopes here due to the plus and minus. What are those slopes?

OpenStudy (anonymous):

trying to figure it out but i can't..

OpenStudy (anonymous):

how about the dy/dx

OpenStudy (wolfe8):

Well, now that you have the equation in terms of y you can find dy/dx. So differentiate that equation. I think you know how to do this part :)

OpenStudy (wolfe8):

Hint: you can rewrite the equation as\[y=\pm \left( 54-2x ^{2} \right)^{1/2}\]

OpenStudy (wolfe8):

@silverxx Should I show you more?

OpenStudy (anonymous):

yes yes. i know..

OpenStudy (anonymous):

wait is -4x/ 2y is -2x/y??

OpenStudy (wolfe8):

Yes

OpenStudy (anonymous):

then help me to get y on this equation please hehehe -2x/y = y-1/ x- 10

OpenStudy (wolfe8):

Hmm. Not sure where I am going myself with this, but if what you meant is \[\frac{ -2x }{ y }=\frac{ y-1 }{ x-10 }\] y is \[\frac{ 1 }{2 }\left( 1-\sqrt{-8x ^{2}+80x+1} \right)\]

OpenStudy (anonymous):

yes yes thats it

OpenStudy (anonymous):

my solution was it turned out to be a quadratic

OpenStudy (anonymous):

dear wolfe8 what program are you using to write your equations so well :)

OpenStudy (phi):

First, I would use implicit differentiation to find dy/dx in 2x^2 + y^2 = 54 4x + 2y dy/dx =0 dy/dx= -2x/y

OpenStudy (phi):

once you have the equation \[ \frac{ -2x }{ y }=\frac{ y-1 }{ x-10 } \] cross multiply to get \[ -2x^2 +20x = y^2 -y\\ y+20x= 2x^2+y^2 \] from the original equation 2x^2 + y^2 = 54 we can simplify \[y+20x= 2x^2+y^2\\y+20x=54 \\ y= 54-20x\] substitute for y in the original equation \[ 2x^2+y^2=54\\ 2x^2 +(54-x)^2=54 \] that gives a hairy quadratic with roots x=3 and x= 159/67

OpenStudy (phi):

to find the tangent lines, we can solve for the slope, and use the point slope form of the equation of a line. first find y, using y = 54-20x. with x =159/67 we get y = 438/67 the slope m= -2x/y = -318/438= - 53/73 knowing (10,1) is on the line, we have y - 1 = (-53/73) (x-10) -73y +73 = -53x +530 53x -73y -603=0 you can follow the same procedure using the other root x=3 to find the equation of the other tangent line.

OpenStudy (anonymous):

phi, how do you type out the equations so clearly, they appear just like in the text books :)

OpenStudy (phi):

use the equation editor (see button on lower left of input area)

OpenStudy (wolfe8):

@carlyn101 Sorry I didn't see your question. As phi said, it is the editor.

OpenStudy (phi):

it is latex so you can type it directly in, using \ [ \ ] (with no spaces between the slash and bracket. or \ ( \ ) for embedded forms. example \( x^2 \) also, right click on an equation and select Show Math as Tex commands if you want to copy and paste some complicated expression

OpenStudy (anonymous):

I just typed in this --> 2x^2 + y^2 = 54 then i put it in the equation editor, then hit the INSERT button and nothing happened ?

OpenStudy (anonymous):

Also do you know if i can do equations on a wacom writing tablet ?

OpenStudy (phi):

@carlyn101 yes, open the equation editor, type into its input area you should see the latex form in the large box above the input area use the buttons on the right side to select formats.. e.g. square root, integral sign, etc when finished with the equation, select the INSERT button in the equation editor. The equation will show up in the "input area" where you type in text. Select POST to publish the result.

OpenStudy (anonymous):

thank you thank you @phi

OpenStudy (phi):

I noticed I lost a sign at the very end... knowing (10,1) is on the line, we have y - 1 = (-53/73) (x-10) -73y +73 = 53x - 530 (multiply both sides by -73, distribute +53) 53x +73y -603=0 now it matches your answer up top.

OpenStudy (anonymous):

i have my answer already.. i didnt give up :) happy having same solution with you ;)

OpenStudy (anonymous):

\[2x^2 + y^2 = 54\] thanks phi and wolf